The average mass of 50 pumpkins is 2,1 kg. If one more pumpkin is added, the average mass is 2,2 kg. What is the mass of the extra pumpkin?


1
Expert's answer
2021-07-21T15:03:20-0400

"total mass=50\u00d72.1=105" kg

"let x" be equal to weight of the added pumpkin.

"=(105+x)\div51=2.2"


"=\frac{(105+x)}{51}×51=2.2×51"


"105+x=112.2"


"x=112.2-105=7.2"

The mass of the extra pumpkin is 7.2 kilograms.